how do i get the lateral area and surface area of a cone where i only have the hieght=8 and slant hieght=10
Answer by venugopalramana(3286)

RADIUS(R),SLANT HEIGHT(L) AND VERTICAL HT.(H)ARE RELATED BY PYTHOGARUS THEOREM AS
L^2=H^2+R^2……..10^2=8^2+R^2
R^2=100-64=36
R=6
LATERAL S.A.=PI*R*L=3.14*6*10=188.4
RADIUS(R),SLANT HEIGHT(L) AND VERTICAL HT.(H)ARE RELATED BY PYTHOGARUS THEOREM AS
L^2=H^2+R^2……..10^2=8^2+R^2
R^2=100-64=36
R=6
LATERAL S.A.=PI*R*L=3.14*6*10=188.4